May 31 Accrued interest revenue for the year. Accrue interest revenue on all of the notes in one entry. (Use a 365-day year fPlease answer the question correctly, and give the progress. Note: 418 is incorrect
Answer #1

Interest revenue = 15000*9%*3/12+14000*6%*25/365+4000*12%*15/365 = 338+58+20 = 416

Journal entry

Date account and explanation Debit Credit
May 31 Interest receivable 416
Interest revenue 416
(To record accrued interest)